OpenStudy (anonymous):

list the intercepts for y=x^4-2x^2-8

OpenStudy (anonymous):

y intercept is -8 you need to solve for x to find x intercepts

OpenStudy (anonymous):

how is y intercept -8?

OpenStudy (anonymous):

you get y intercept when x=0

OpenStudy (anonymous):

\[y=x ^{4}-2x ^{2}-8\] list the intercepts

OpenStudy (dumbcow):

you can factor this notice -4*2 = -8 -4 +2 = -2 (x^2 +2)(x^2 -4) = 0 now solve each factor separately x^2 -4 = 0 x^2 = 4 x = +-(2) x^2 +2 = 0 x^2 = -2 x = +-sqrt(2)i

OpenStudy (anonymous):

plug x=0 into equation: y = 0 + 0 - 8

OpenStudy (anonymous):

oohh i get it.. thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if you are happy with the answer plz close the question by clicking on good answer

OpenStudy (dumbcow):

yes